haleydawkins:

Darius is studying the relationship between mathematics and art. He asks friends to each draw a "typical” rectangle. He measures the length and width in centimeters of each rectangle and plots the points on a graph, where x represents the width and y represents the length. The points representing the rectangles are (6.1, 12.0), (5.0, 8.1), (9.1, 15.2), (6.5, 10.2), (7.4, 11.3), and (10.9, 17.5). Which equation could Darius use to determine the length, in centimeters, of a "typical” rectangle for a given width in centimeters?

Mercury:

choices: A. y = 0.605x + 0.004 B. y = 0.959x + 0.041 C. y = 1.518x + 0.995 D. y = 1.967x + 0.984 he wants a predictive equation that will give a predicted value of y for a given x value, based on the data he's already collected. we should find a least-squares regression equation. you *could* do this by hand but it's very complicated and much easier to plug your data into a least squares regression line calculator https://www.socscistatistics.com/tests/regression/default.aspx